OpenStudy (anonymous):

integral 1/(1-tan^2x)

OpenStudy (anonymous):

\[\int\limits_{}^{}1/(1-\tan^2x)\]

OpenStudy (anonymous):

I think we need to do u = 1-tan^2x I'm trying that now

OpenStudy (solomonzelman):

\(\large\color{slate}{\displaystyle\int\limits_{~}^{~}\frac{1}{1-\tan^2{x}}~dx}\) \(\large\color{slate}{\displaystyle\int\limits_{~}^{~}\frac{1}{\dfrac{\cos^2x}{\cos^2x}-\dfrac{\sin^2x}{\cos^2{x}}}~dx}\) \(\large\color{slate}{\displaystyle\int\limits_{~}^{~}\frac{1}{\dfrac{\cos^2x-\sin^2x}{\cos^2{x}}}~dx}\) \(\large\color{slate}{\displaystyle\int\limits_{~}^{~}\frac{1\times \cos^2x}{\dfrac{\cos^2x-\sin^2x}{\cos^2{x}}\times \cos^2x}~dx}\) \(\large\color{slate}{\displaystyle\int\limits_{~}^{~}\frac{\cos^2x}{\cos^2x-\sin^2{x}}~dx}\)

OpenStudy (solomonzelman):

no it is a mess my brain is not working right now

OpenStudy (anonymous):

It's ok this problem is also making my brain hurt lol

OpenStudy (empty):

Hmmm is it being the difference of two squares helpful at all?

OpenStudy (anonymous):

How about cos2x = cos^2x - sin^2x And cos^2x= cos2x+sin^2x

OpenStudy (solomonzelman):

Empty i thought of that, but what would u then do

OpenStudy (anonymous):

\[\int\limits_{}^{}(\cos2x + \sin^2x)/\cos2x\]

OpenStudy (anonymous):

Integral of 1 + integral of sin^2x/cos2x

OpenStudy (solomonzelman):

cos²x ?

OpenStudy (solomonzelman):

or cos(2x) ?

OpenStudy (anonymous):

I'm confused which cos you're talking about

OpenStudy (solomonzelman):

both

OpenStudy (anonymous):

cos2x = cos^2x - sin^2x I used it to change cos^2x/(cos^2x - sin^2x)

OpenStudy (anonymous):

I can't see what to do if we do use the difference of squares

OpenStudy (empty):

that was a glitch it double posted idk why lol

OpenStudy (solomonzelman):

yeah difference of squares doesn't get me anywhere, then what sub would I make? and go into world's worse du=....

OpenStudy (anonymous):

I used wolfram to solve it and based on the answer there should be a logarithm using u substitution

OpenStudy (anonymous):

And a 1 somewhere

OpenStudy (anonymous):

oh wait. If at the beginning, we do u = tanx Then its du = sec^2xdx du=(1+tan^2x)dx So it would be \[\int\limits_{}^{}1/((1-u^2)(1+u^2))\]

OpenStudy (anonymous):

And that would be solvable with partial fractions I think

OpenStudy (solomonzelman):

no master

OpenStudy (solomonzelman):

sorry to dissapoint you, but your dx = (will be replaced by) cos²(idk what)

OpenStudy (solomonzelman):

lol, du=sec^2x dx cos^2x du=dx cos^2(tan\(^{-1}u\)) du=dx

OpenStudy (anonymous):

u = tanx du=sec^2xdx cos^2xdu=dx So \[\int\limits_{}^{}\cos^2x/(1-u^2)du\]

OpenStudy (solomonzelman):

can not have that-:( x's with u' and du like that

OpenStudy (anonymous):

But can't we just put the cos^2x as sec^2x in the denominator. And then sec^2x = 1+tan^2x = 1+u^2

OpenStudy (anonymous):

Even if you substitute it back? I didn't know

OpenStudy (solomonzelman):

u= tan(x) \(\large\color{slate}{\displaystyle\int\limits_{~}^{~}\frac{\cos^2x}{1-u^2}~du}\) \(\large\color{slate}{\displaystyle\int\limits_{~}^{~}\frac{1}{(1+\tan^2x)(1-u^2)}~du}\) \(\large\color{slate}{\displaystyle\int\limits_{~}^{~}\frac{1}{(1+u^2)(1-u^2)}~du}\)

OpenStudy (solomonzelman):

yes, sounds like it works.... no partial fractions is the intitive.

OpenStudy (solomonzelman):

good job, i should say, master. indeed

OpenStudy (solomonzelman):

now, partial fractions*

OpenStudy (anonymous):

Thank you lol

OpenStudy (solomonzelman):

\(\large\color{slate}{\displaystyle\int\limits_{~}^{~}\frac{1}{(1+\tan^2x)(1-u^2)}~du}\) \(\large\color{slate}{\displaystyle\frac{1}{4}\int\limits_{~}^{~}\frac{2}{u^2+1}+\frac{1}{u+1}-\frac{1}{u-1}du}\)

OpenStudy (solomonzelman):

oh, the tan^2x is u^2

OpenStudy (solomonzelman):

so if that substitution actually works as we said, the rest should be easy.

OpenStudy (anonymous):

Yup that matches what wolfram said the answer was.

OpenStudy (anonymous):

Good job and thank you :)

OpenStudy (solomonzelman):

Lol, I need to get back to one variable integration... this is a good exercise.

OpenStudy (solomonzelman):

I had another idea, that I messaged to Empti

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle \frac{1}{1-x} =\sum_{n=0}^{\infty}x^n }\) \(\large\color{black}{ \displaystyle \frac{1}{1-\tan(x)} =\sum_{n=0}^{\infty}\tan^n(x) }\) \(\large\color{black}{ \displaystyle \frac{1}{1-\tan^2(x)} =\sum_{n=0}^{\infty}\tan^{2n}(x) }\) \(\large\color{black}{ \displaystyle \int\frac{1}{1-\tan^2(x)} dx=\int \sum_{n=0}^{\infty}\tan^{2n}(x)dx }\)

OpenStudy (solomonzelman):

oh, no then, this becomes a difficult integral on the right

OpenStudy (solomonzelman):

oh, you will learn it soon then.... it is just a power series representation thing very easy. good luck

OpenStudy (anonymous):

Thanks and to you as well

OpenStudy (solomonzelman):

TH\(\color{blue}{\unicode{x13D49}}\)NK \(\color{blue}{\unicode{x3D4}}\)\(\color{blue}{O}\)U

OpenStudy (solomonzelman):

cu